You have 1.25 L of a 2.00 M solution of Hypochlorous acid, HClO to which you will add 1.5 M NaOH to adjust the pH.

a) what is the pH of hte solution when 0.500 L of NaOH is added?

b) How many mL of soldium hydroxide should be added to prepare the 7.00 pH buffer solution?

c) If 75.0 mL of 1.5 M Ba(OH)2 is added to the buffer in part 'b' when it is starting at pH 7.00 before adding Ba(OH)2, what is the new pH? Ba(OH)2 is a strong dihydroxic base.
Answer #1

1.25 L of 2.oom hyper chlorous acid = 1.25XX 2.00 ml of hyperchlorousund. = 2.5 mol of hyperchlorous aced. @ 0.57 ofis m nuo0.75/1.75 thus from eae we have pH = 7.53 + log - 1711-7- = 7.53 +log at 57-16 () ut, v rool of Nach should be added. me of 1(C) Bacon) ----> Balt+- 2014 - 1.5m 2x1.5m Bacom - -1.5m : LOH) = 2x1m = 3m - 0.385x1.5ml = 0.5775 mok. from the port (6) til